OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = (x-7)/(x+2) and g(x) = (-2x-7)/(x-1) Need step by step thank you:)

OpenStudy (anonymous):

Try plugging \(g(x)\) in as \(x\) in the definition of \(f(x)\) and see if it simplifies to x.

OpenStudy (anonymous):

so f((-2x-7)/(x-1))

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

uh x = -2 - 9/x-1?

OpenStudy (anonymous):

You need to simplify this: \[\frac{-\frac{2x + 7}{x - 1} - 7}{-\frac{2x + 7}{x - 1} + 2}\]